### 01 Dec 2023 v2

@@ 1 Introduction
In this paper, we are concerned with a particular subclass of (CPOP), where the polynomials $f,g_{1},\ldots,g_{m},h_{1},\ldots,h_{l}$ all have real coefficients, or equivalently, the polynomials $f,g_{1},\ldots,g_{m},h_{1},\ldots,h_{l}$ are invariant under the conjugate transformation: ${\mathbf{z}}\mapsto\overline{{\mathbf{z}}}$. For this type of (CPOP), we propose a new moment-HSOS hierarchy that involves positive semidefinite real matrices. We show that, in comparison to the complex moment-HSOS hierarchy, this new hierarchy (1) has lower complexity (see Table 1), and (2) provides the same bound at each relaxation order (Theorem 3.5). Furthermore, as another contribution of this paper, we prove that global optimality of the new hierarchy is attained when the ranks of the moment matrix and certain submatrix equal two in the presence of a sphere constraint. Note that here we do not require the “joint hyponormality” condition to declare global optimality (cf. [12, Proposition 4.1]). The rank-two condition also implies that the related complex polynomial optimization problem has either two real optimal solutions or a pair of conjugate optimal solutions. In the latter case, we provide a simple procedure for extracting a pair of conjugate global minimizers from the moment matrix of order one.
The efficiency of the new hierarchy is demonstrated on randomly generated CPOP instances as well as two CPOPs (respectively related to Smale’s Mean Value conjecture and the Mordell inequality conjecture) from the literature. Finally, as an application, we apply the new hierarchy to the problem of polyphase code design arising from signal processing.
@@ 2 Notation and preliminaries
Let ${\mathbb{N}}$ be the set of nonnegative integers. For $n\in{\mathbb{N}}\setminus\{0\}$, let $[n]\coloneqq\{1,2,\ldots,n\}$. For ${\boldsymbol{\alpha}}=(\alpha_{i})\in{\mathbb{N}}^{n}$, let $|{\boldsymbol{\alpha}}|\coloneqq\sum_{i=1}^{n}\alpha_{i}$. For $r\in{\mathbb{N}}$, let ${\mathbb{N}}^{n}_{r}\coloneqq\{{\boldsymbol{\alpha}}\in{\mathbb{N}}^{n}\mid|{\boldsymbol{\alpha}}|\leq r\}$. We use $A\succeq 0$ to indicate that the matrix $A$ is positive semidefinite (PSD). Let i be the imaginary unit, satisfying $\hbox{\bf{i}}^{2}=-1$. Let ${\mathbf{z}}=(z_{1},\ldots,z_{n})$ be a tuple of complex variables and $\overline{{\mathbf{z}}}=(\overline{z}_{1},\ldots,\overline{z}_{n})$ its conjugate. Let $\overline{a}$ denote the conjugate of a complex number $a$ and ${\boldsymbol{v}}^{*}$ (resp. $A^{*}$) denote the conjugate transpose of a complex vector ${\boldsymbol{v}}$ (resp. a complex matrix $A$). We denote by ${\mathbb{C}}[{\mathbf{z}},\overline{{\mathbf{z}}}]\coloneqq{\mathbb{C}}[z_{1},\ldots,z_{n},\overline{z}_{1},\ldots,\overline{z}_{n}]$, ${\mathbb{R}}[{\mathbf{z}},\overline{{\mathbf{z}}}]\coloneqq{\mathbb{R}}[z_{1},\ldots,z_{n},\overline{z}_{1},\ldots,\overline{z}_{n}]$, the complex and real polynomial rings in ${\mathbf{z}},\overline{{\mathbf{z}}}$, respectively. A polynomial $f\in{\mathbb{C}}[{\mathbf{z}},\overline{{\mathbf{z}}}]$ (resp. ${\mathbb{R}}[{\mathbf{z}},\overline{{\mathbf{z}}}]$) can be written as $f=\sum_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{n}\times{\mathbb{N}}^{n}}f_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}{\mathbf{z}}^{{\boldsymbol{\beta}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}}$ with $f_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\in{\mathbb{C}}$ (resp. $f_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\in{\mathbb{R}}$), ${\mathbf{z}}^{{\boldsymbol{\beta}}}=z_{1}^{\beta_{1}}\cdots z_{n}^{\beta_{n}},\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}}=\overline{z}_{1}^{\gamma_{1}}\cdots\overline{z}_{n}^{\gamma_{n}}$. The conjugate of $f$ is $\overline{f}=\sum_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{n}\times{\mathbb{N}}^{n}}\overline{f}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}{\mathbf{z}}^{{\boldsymbol{\gamma}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\beta}}}$. The polynomial $f$ is self-conjugate if $\overline{f}=f$. It is clear that self-conjugate polynomials take only real values. The set of self-conjugate polynomials in ${\mathbb{C}}[{\mathbf{z}},\overline{{\mathbf{z}}}]$ (resp. ${\mathbb{R}}[{\mathbf{z}},\overline{{\mathbf{z}}}]$) is denoted by ${\mathbb{C}}[{\mathbf{z}},\overline{{\mathbf{z}}}]^{\rm{c}}$ (resp. ${\mathbb{R}}[{\mathbf{z}},\overline{{\mathbf{z}}}]^{\rm{c}}$). For $r\in{\mathbb{N}}$, let us denote
@@ 4 Detecting global optimality and extracting solutions
In general, detecting global optimality and extracting solutions for (RMom-$r$) can be carried out in the same way as for (CMom-$r$). For details, we refer the reader to Proposition 4.1 of [12] and the discussions therein. However, we observe that the ‘joint hyponormality’ condition in [12, Proposition 4.1] could be actually replaced with the following simpler ‘hyponormality’ condition.
@@
The proof is almost the same as the one of [12, Proposition 4.1], which relies on [12, Theorem 5.1]. One just needs to note that the condition (Hyponormality) implies that the so-called shift operators are all normal, which together with pairwise commutativity suffices to guarantee that these shift operators (and their adjoint operators) are simultaneously diagonalizable.
@@
As a result, $A_{1},\ldots,A_{n}$ are simultaneously diagonalizable.
@@
The proof is an adaptation of the one of [12, Theorem 5.1].
@@
Then by Proposition 4.3, $T_{1},\ldots,T_{n}$ are simultaneously diagonalizable. That is, there exists a unitary matrix $P$ such that $T_{i}=PD_{i}P^{*}$, $i=1,\ldots,n$, where $D_{i}=\operatorname{diag}\,(d_{i1},d_{i2})$ is a diagonal matrix. For convenience, we let $T^{{\boldsymbol{\alpha}}}\coloneqq T_{1}^{\alpha_{1}}\cdots T_{n}^{\alpha_{n}}$. Then for $|{\boldsymbol{\beta}}|,|{\boldsymbol{\gamma}}|\leq r$, we have
@@
The sphere constraint in Theorem 4.5 can be replaced by the unit-norm constraint on each variable as the latter together with $\hbox{\rm{rank}}\,{{\mathbf{M}}_{r}({\mathbf{y}})}=\hbox{\rm{rank}}\,{{\mathbf{M}}_{r-2}({\mathbf{y}})}$ can also ensure that $T_{1},\ldots,T_{n}$ are simultaneously diagonalizable; see [12, Appendix B].
@@
Thus, we can retrieve the pair of conjugate optimal solutions from the above ‘Cholesky decomposition’ of ${\mathbf{M}}_{1}({\mathbf{y}})$.
@@ 5 Numerical experiments
The real moment-HSOS hierarchy has been implemented in the Julia package TSSOS [^3] TSSOS is freely available at https://github.com/wangjie212/TSSOS.. In this section, we compare the performance of the real moment-HSOS hierarchy (R-HSOS), the complex moment-HSOS hierarchy (C-HSOS), and the real moment-SOS hierarchy (R-SOS) on CPOPs with real coefficients using TSSOS where Mosek 10.0 [2] is employed as an SDP solver with default settings. In presenting the results, the column labelled by ‘opt’ records optima of SDPs and the column labelled by ‘time’ records running time in seconds. Moreover, the symbol ‘-’ means that Mosek runs out of memory.
@@
where $[{\mathbf{z}}]_{1}=({\mathbf{z}}^{{\boldsymbol{\alpha}}})_{|{\boldsymbol{\alpha}}|\leq 1}$ and $Q$ is a random symmetric matrix whose entries are selected with respect to the uniform probability distribution on $[-1,1]$.
We approach (15) for $n=50,100,\ldots,300$ with the first-order relaxation of the three hierarchies. To measure the quality of obtained lower bounds, we also compute an upper bound (indicated by ’ub’ in the table) for (15) using the nonlinear programming solver Ipopt. Related results are shown in Table 2, from which we see that R-HSOS is the fastest and is more scalable with the size of the problem. Interestingly, we can also see that the lower bounds produced by the SDP relaxations coincide with the corresponding upper bounds, and thus are actually globally optimal.
@@
where $[{\mathbf{z}}]_{2}=({\mathbf{z}}^{{\boldsymbol{\alpha}}})_{|{\boldsymbol{\alpha}}|\leq 2}$ and $Q$ is a random symmetric matrix whose entries are selected with respect to the uniform probability distribution on $[-1,1]$.
We approach (16) for $n=8,10,\ldots,22$ with the second-order relaxations of the three hierarchies. To measure the quality of obtained lower bounds, we also compute an upper bound (indicated by ’ub’ in the table) for (16) using Ipopt. Related results are shown in Table 3, from which we see that R-HSOS is the fastest and is more scalable with the size of the problem. Besides, we can see that the lower bounds produced by R-SOS coincide with the corresponding upper bounds, and thus are globally optimal, while the lower bounds produced by R-HSOS and C-HSOS are not globally optimal.
@@
Theorem 5.1 was proved by Smale in 1981, and he further asked whether one can replace the factor $4$ in (17) by 1, or even possibly by $\frac{n}{n+1}$, which is known as Smale’s Mean Value conjecture and also appears in [23]. This conjecture is verified for $n=1,2,3$ [20, Chapter 7.2] and has the following equivalent (normalized) form.
@@
We approach (20) for $n=2,3,4,5$ with R-HSOS and C-HSOS. Related results are shown in Table 4. For $n=2$, the $2$nd order relaxation (numerically) retrieves the desired value $\frac{2}{3}$; for $n=3$, the $5$th order relaxation (numerically) retrieves the desired value $\frac{3}{4}$; for $n=4$, the $9$th order relaxation (numerically) retrieves the desired value $\frac{4}{5}$; for $n=5$, the hierarchies are not able to retrieve the desired value $\frac{5}{6}$ due to memory limitation. It can be seen from the table that R-HSOS is more scalable with the size of the problem.
@@
$n$ $r$ R-HSOS C-HSOS
opt time opt time
2 2 0.66666 0.005 0.66666 0.006
3 5 0.75000 0.15 0.75000 0.45
4 9 0.80000 6257 - -
5 7 1.20615 16497 - -
@@
We approach (23) with R-HSOS and C-HSOS. Related results are shown in Table 5 ($n=3$) and Table 6 ($n=4$), from which we see that the upper bounds provided by the two hierarchies get close to the desired value $27$ or $256$ as the relaxation order $r$ grows while R-HSOS is faster and more scalable.

### 25 Feb 2024 v3

@@ 1 Introduction
In this paper, we are concerned with a particular subclass of (CPOP), where the polynomials $f,g_{1},\ldots,g_{m},h_{1},\ldots,h_{l}$ all have real coefficients, or equivalently, the polynomials $f,g_{1},\ldots,g_{m},h_{1},\ldots,h_{l}$ are invariant under the conjugate transformation: ${\mathbf{z}}\mapsto\overline{{\mathbf{z}}}$. Many complex polynomial optimization problems involving norms (e.g., the problem of polyphase code design) belong to this class (see Sections 5 and 6). For this type of (CPOP), we propose a symmetry-adapted moment-HSOS hierarchy that involves positive semidefinite real matrices. We show that, in comparison to the complex moment-HSOS hierarchy, this new hierarchy (1) has lower complexity (see Table 1), and (2) provides the same bound at each relaxation order (Theorem 3.5). Essentially, this result is an extension of Riener et al.’s work [23] to complex polynomial optimization with conjugate symmetry. Furthermore, as another contribution of this paper, we prove that global optimality of the new hierarchy is attained when the ranks of the moment matrix and certain submatrix equal two in the presence of a sphere constraint. Note that here we do not require the “joint hyponormality” condition to declare global optimality (cf. [14, Proposition 4.1]). The rank-two condition also implies that the related complex polynomial optimization problem has either two real optimal solutions or a pair of conjugate optimal solutions. In the latter case, we provide a simple procedure for extracting a pair of conjugate global minimizers from the moment matrix of order one.
The efficiency of the new hierarchy is demonstrated on randomly generated CPOP instances as well as two CPOPs (respectively related to Smale’s Mean Value conjecture and the Mordell inequality conjecture) from the literature. Finally, as an application, we apply the new hierarchy to the important problem of polyphase code design arising from signal processing.
@@ 2 Notation and preliminaries
Let ${\mathbb{N}}$ be the set of nonnegative integers. For $n\in{\mathbb{N}}\setminus\{0\}$, let $[n]\coloneqq\{1,2,\ldots,n\}$. For ${\boldsymbol{\alpha}}=(\alpha_{i})\in{\mathbb{N}}^{n}$, let $|{\boldsymbol{\alpha}}|\coloneqq\sum_{i=1}^{n}\alpha_{i}$. For $r\in{\mathbb{N}}$, let ${\mathbb{N}}^{n}_{r}\coloneqq\{{\boldsymbol{\alpha}}\in{\mathbb{N}}^{n}\mid|{\boldsymbol{\alpha}}|\leq r\}$. We use $A\succeq 0$ to indicate that the matrix $A$ is positive semidefinite (PSD). Let i be the imaginary unit, satisfying $\hbox{\bf{i}}^{2}=-1$. Let ${\mathbf{z}}=(z_{1},\ldots,z_{n})$ be a tuple of complex variables and $\overline{{\mathbf{z}}}=(\overline{z}_{1},\ldots,\overline{z}_{n})$ its conjugate. Let $\overline{a}$ denote the conjugate of a complex number $a$ and ${\boldsymbol{v}}^{*}$ (resp. $A^{*}$) denote the conjugate transpose of a complex vector ${\boldsymbol{v}}$ (resp. a complex matrix $A$). We denote by ${\mathbb{C}}[{\mathbf{z}},\overline{{\mathbf{z}}}]\coloneqq{\mathbb{C}}[z_{1},\ldots,z_{n},\overline{z}_{1},\ldots,\overline{z}_{n}]$, ${\mathbb{R}}[{\mathbf{z}},\overline{{\mathbf{z}}}]\coloneqq{\mathbb{R}}[z_{1},\ldots,z_{n},\overline{z}_{1},\ldots,\overline{z}_{n}]$, the complex and real polynomial rings in ${\mathbf{z}},\overline{{\mathbf{z}}}$, respectively. A polynomial $f\in{\mathbb{C}}[{\mathbf{z}},\overline{{\mathbf{z}}}]$ (resp. ${\mathbb{R}}[{\mathbf{z}},\overline{{\mathbf{z}}}]$) can be written as $f=\sum_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{n}\times{\mathbb{N}}^{n}}f_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}{\mathbf{z}}^{{\boldsymbol{\beta}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}}$ with $f_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\in{\mathbb{C}}$ (resp. $f_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}\in{\mathbb{R}}$), ${\mathbf{z}}^{{\boldsymbol{\beta}}}=z_{1}^{\beta_{1}}\cdots z_{n}^{\beta_{n}},\overline{{\mathbf{z}}}^{{\boldsymbol{\gamma}}}=\overline{z}_{1}^{\gamma_{1}}\cdots\overline{z}_{n}^{\gamma_{n}}$. The conjugate of $f$ is $\overline{f}=\sum_{({\boldsymbol{\beta}},{\boldsymbol{\gamma}})\in{\mathbb{N}}^{n}\times{\mathbb{N}}^{n}}\overline{f}_{{\boldsymbol{\beta}},{\boldsymbol{\gamma}}}{\mathbf{z}}^{{\boldsymbol{\gamma}}}\overline{{\mathbf{z}}}^{{\boldsymbol{\beta}}}$. The polynomial $f$ is self-conjugate if $\overline{f}=f$. It is clear that self-conjugate polynomials take only real values. The set of self-conjugate polynomials in ${\mathbb{C}}[{\mathbf{z}},\overline{{\mathbf{z}}}]$ (resp. ${\mathbb{R}}[{\mathbf{z}},\overline{{\mathbf{z}}}]$) is denoted by ${\mathbb{C}}[{\mathbf{z}},\overline{{\mathbf{z}}}]^{\rm{c}}$ (resp. ${\mathbb{R}}[{\mathbf{z}},\overline{{\mathbf{z}}}]^{\rm{c}}$). For example, we have $3-|z_{1}|^{2}+\frac{1+\hbox{\bf{i}}}{2}z_{1}\overline{z}_{2}^{2}+\frac{1-\hbox{\bf{i}}}{2}z_{2}^{2}\overline{z}_{1}\in{\mathbb{C}}[{\mathbf{z}},\overline{{\mathbf{z}}}]^{\rm{c}}$ and $3-|z_{1}|^{2}+\frac{1}{2}z_{1}\overline{z}_{2}^{2}+\frac{1}{2}z_{2}^{2}\overline{z}_{1}\in{\mathbb{R}}[{\mathbf{z}},\overline{{\mathbf{z}}}]^{\rm{c}}$. For $r\in{\mathbb{N}}$, let us denote
@@ 4 Detecting global optimality and extracting solutions
In general, detecting global optimality and extracting solutions for (RMom-$r$) can be carried out in the same way as for (CMom-$r$). For details, we refer the reader to Proposition 4.1 of [14] and the discussions therein. However, we observe that the “joint hyponormality” condition in [14, Proposition 4.1] could be actually replaced with the following simpler “hyponormality” condition.
@@
The proof is almost the same as the one of [14, Proposition 4.1], which relies on [14, Theorem 5.1]. One just needs to note that the condition (Hyponormality) implies that the so-called shift operators are all normal, which together with pairwise commutativity suffices to guarantee that these shift operators (and their adjoint operators) are simultaneously unitarily diagonalizable; see Theorem 2.5.5 of [13] (and Theorem 1 of [11]).
@@
As a result, $A_{1},\ldots,A_{n}$ are simultaneously unitarily diagonalizable.
@@
The proof is an adaptation of the one of [14, Theorem 5.1].
@@
Then by Proposition 4.3, $T_{1},\ldots,T_{n}$ are simultaneously unitarily diagonalizable. That is, there exists a unitary matrix $P$ such that $T_{i}=PD_{i}P^{*}$, $i=1,\ldots,n$, where $D_{i}=\operatorname{diag}\,(d_{i1},d_{i2})$ is a diagonal matrix. For convenience, we let $T^{{\boldsymbol{\alpha}}}\coloneqq T_{1}^{\alpha_{1}}\cdots T_{n}^{\alpha_{n}}$. Then for $|{\boldsymbol{\beta}}|,|{\boldsymbol{\gamma}}|\leq r$, we have
@@
The sphere constraint in Theorem 4.5 can be replaced by the unit-norm constraint on each variable as the latter together with $\hbox{\rm{rank}}\,{{\mathbf{M}}_{r}({\mathbf{y}})}=\hbox{\rm{rank}}\,{{\mathbf{M}}_{r-2}({\mathbf{y}})}$ can also ensure that $T_{1},\ldots,T_{n}$ are simultaneously unitarily diagonalizable; see [14, Appendix B].
@@
Thus, we can retrieve the pair of conjugate optimal solutions from the above “Cholesky decomposition” of ${\mathbf{M}}_{1}({\mathbf{y}})$.
@@ 5 Numerical experiments
The real moment-HSOS hierarchy has been implemented in the Julia package TSSOS [^3] TSSOS is freely available at https://github.com/wangjie212/TSSOS.. In this section, we compare the performance of the real moment-HSOS hierarchy (R-HSOS), the complex moment-HSOS hierarchy (C-HSOS), and the real moment-SOS hierarchy (R-SOS) on CPOPs with real coefficients using TSSOS where Mosek 10.0 [2] is employed as an SDP solver with default settings. For R-SOS, we have exploited sign symmetry to reduce the SDP size [32]. In presenting the results, the column labelled by ‘opt’ records optima of SDPs and the column labelled by ‘time’ records running time in seconds. Moreover, the symbol ‘-’ means that Mosek runs out of memory.
@@
where $[{\mathbf{z}}]_{1}=({\mathbf{z}}^{{\boldsymbol{\alpha}}})_{|{\boldsymbol{\alpha}}|\leq 1}$ and $Q$ is a random real symmetric matrix whose entries are selected with respect to the uniform probability distribution on $[-1,1]$.
We approach (15) for $n=50,100,\ldots,300$ with the first-order relaxation of the three hierarchies. To measure the quality of obtained lower bounds, we also compute an upper bound (indicated by ‘ub’ in the table) for (15) using the nonlinear programming solver Ipopt. The related results are shown in Table 2, from which we see that R-HSOS is the fastest and is more scalable with the size of the problem. Interestingly, we can also see that the lower bounds produced by the SDP relaxations coincide with the corresponding upper bounds, and thus are actually globally optimal.
@@
where $[{\mathbf{z}}]_{2}=({\mathbf{z}}^{{\boldsymbol{\alpha}}})_{|{\boldsymbol{\alpha}}|\leq 2}$ and $Q$ is a random real symmetric matrix whose entries are selected with respect to the uniform probability distribution on $[-1,1]$.
We approach (16) for $n=8,10,\ldots,22$ with the second-order relaxations of the three hierarchies. To measure the quality of obtained lower bounds, we also compute an upper bound (indicated by ‘ub’ in the table) for (16) using Ipopt. The related results are shown in Table 3, from which we see that R-HSOS is the fastest and is more scalable with the size of the problem. Besides, we can see that the lower bounds produced by R-SOS coincide with the corresponding upper bounds, and thus are globally optimal, while the lower bounds produced by R-HSOS and C-HSOS are not globally optimal.
@@
Theorem 5.1 was proved by Smale in 1981, and he further asked whether one can replace the factor $4$ in (17) by 1, or even possibly by $\frac{n}{n+1}$, which is known as Smale’s Mean Value conjecture and also appears in [25]. This conjecture is verified for $n=1,2,3$ [22, Chapter 7.2] and has the following equivalent (normalized) form.
@@
We approach (20) for $n=2,3,4,5$ with R-HSOS, C-HSOS, and R-SOS. To measure the quality of obtained upper bounds, we also compute an lower bound (indicated by ‘lb’ in the table) for (20) using Ipopt. The related results are shown in Table 4. For $n=2$, the $2$nd order relaxation (numerically) retrieves the desired value $\frac{2}{3}$; for $n=3$, the $5$th order relaxation (numerically) retrieves the desired value $\frac{3}{4}$; for $n=4$, the $9$th order relaxation (numerically) retrieves the desired value $\frac{4}{5}$; for $n=5$, the hierarchies are not able to retrieve the desired value $\frac{5}{6}$ due to memory limitation. It can be seen from the table that R-HSOS is more scalable with the size of the problem. Note that for $n\geq 4$, the local solver fails in returning lower bounds.
@@
$n$ $r$ lb R-HSOS C-HSOS R-SOS
opt time opt time opt time
2 2 0.66666 0.66666 0.005 0.66666 0.006 0.66666 0.009
3 5 0.75000 0.75000 0.15 0.75000 0.45 0.75000 36.6
4 9 *** 0.80000 6257 - - - -
5 7 *** 1.20615 16497 - - - -
@@
We approach (23) with R-HSOS and C-HSOS. The related results are shown in Table 5 ($n=3$) and Table 6 ($n=4$), from which we see that the upper bounds provided by the two hierarchies get close to the desired value $27$ or $256$ as the relaxation order $r$ grows while R-HSOS is more efficient than C-HSOS. In addition, the local solver and R-SOS with minimum relaxation order return $27$ and $256$ for $n=3,4$, respectively.